OpenStudy (anonymous):

Describe the relationship between the mathematics of calculating simple interest, and the mathematics behind analysing arithmetic sequences. How might simple interest be described as a problem of applied arithmetic sequences?

OpenStudy (abb0t):

Formula for simple interest is Y = A(1 + rX), where Y = cumulative balance, A = original balance, X = number of periods of interest and r = interest rate. Y = A(1 + rX) is a linear function. And basic interest grows linearly, making this an example of an arithmetic sequence.

OpenStudy (anonymous):

Thank you so much!